What will the volume of a sample be if 335 mL of an ideal gas at 24 °C and 1.08 atm is cooled to 17 °C and 0.976 atm?



1
Expert's answer
2022-05-02T05:18:04-0400

P1 = 1.08 atm

V1 = 335 mL

T1 = 24°C + 273.15 = 297.15 K

P2 = 0.976 atm

V2 = ???

T2 = 17°C + 273.15 = 290.15 K


Solution:

Mathematical expression for the combined gas law:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (1.08 atm × 335 mL × 290.15 K) / (0.976 atm × 297.15 K) = 361.964 mL = 362 mL

V2 = 362 mL


Answer: The volume of a sample will be 362 mL